OpenStudy (callisto):

Prove \[\frac{sinx}{x} > \frac{2}{\pi} \ \ , \forall x \in (0, \frac{ \pi }{ 2 })\] How to start?

OpenStudy (bahrom7893):

For some reason I'm thinking limits but it's probably because of Sinx / x and has nothing to do with this

OpenStudy (callisto):

I think we need to take limit.. but... we have something to do before taking limit?

OpenStudy (bahrom7893):

Integrate from 0 to pi/2? And something about the area must tell you something? Idk, lol just rambling on.

OpenStudy (callisto):

Well, I don't think we need integration here.

OpenStudy (callisto):

It's because the prof. hasn't taught integration in the lessons yet (even though I have learnt integration before)

OpenStudy (bahrom7893):

Ohh Ok. There are some proof of function being greater than the other.. hmm

OpenStudy (callisto):

1. it's a decreasing function. 2. \[\lim_{x \rightarrow 0} \frac{sinx}{x}=1\] 3. \[\lim_{x \rightarrow \frac{\pi}{2}} \frac{sin\frac{\pi }{2}}{\frac{\pi}{2}}=\frac{2}{\pi}\] Uh-oh! I'm not on the right track :|

OpenStudy (anonymous):

Why can't you just show that it is uniformly decreasing over the interval and calculate the least value? (Once you have shown that the limit is 1 at start)

OpenStudy (callisto):

Well, that's the problem.. How are you going to show it? Finding its derivative?

OpenStudy (anonymous):

you cannot just say that the roots are at n*pi?

OpenStudy (callisto):

How so?

OpenStudy (anonymous):

You want a proof of that as well...

OpenStudy (callisto):

Just a minute \[\frac{d}{dx}\frac{sinx}{x} = \frac{xcosx-sinx}{x^2}\]If I put it =0, I'll get xcosx - sinx = 0 That's where I stuck at. And I haven't learnt the general solution (hmm, for that n*pi) yet, so I would like to know how you get that.

OpenStudy (anonymous):

I mean x= Tan x is a pain.....

OpenStudy (callisto):

Yes!

OpenStudy (anonymous):

You just have do a graph y=x and Tan x and look or use a CAS

OpenStudy (callisto):

Graphing... is not that good...

OpenStudy (callisto):

If graphing works, why not just graph the function sinx / x to show it's a decreasing function?

OpenStudy (anonymous):

Well, quite...:-)

OpenStudy (anonymous):

What are you doing, Fourier?

OpenStudy (callisto):

No.. limit and derivative.

OpenStudy (anonymous):

That's a bit sneaky, hitting you with sinc function...

OpenStudy (anonymous):

Let's see , all you have to do is show that the first positive root is at pi....

OpenStudy (anonymous):

Sin pi/pi = 0

OpenStudy (anonymous):

And anywhere from 0 to pi is not 0

OpenStudy (anonymous):

That's good enough to show the first root is at pi, right?

OpenStudy (callisto):

Oh... so you're solving sinx/x=0? If so, yes!

OpenStudy (anonymous):

So now you are done, I think (assuming usual smoothness principles).

OpenStudy (turingtest):

I don't see the problem by saying that the limit at 0 is 1>pi/2, and that sinx/x is uniformly decreasing over (0,pi/2), and the limit approaching pi/2=2/pi that shows that sinx/x>2/pi in that region

OpenStudy (callisto):

So, step 1, 2, 3 are correct. But how to show step 1 is the point. I'm not quite sure if I can get estudier's idea..

OpenStudy (turingtest):

how to show it's decreasing? find critical points\[f'=\frac{x\cos x-\sin x}{x^2}=0\implies x\cot x=0\]so the critical points are\[x=\{0,\frac\pi2\}\]hence you just need to test a point in that region, say \(x=\pi/4\), and that tells you whether the function is increasing or decreasing on that interval

OpenStudy (callisto):

Wait.. how.. why... xcosx - sinx = 0 xcotx -1 = 0 xcotx = 1 Am I doing anything wrong here?

OpenStudy (turingtest):

oh shnap I am out to lunch, but I still say we find that critical point and just test the interval let me think how....

OpenStudy (callisto):

Don't think while you eat!

OpenStudy (turingtest):

I know I have seen a proof that tanx>x in (0,pi/2) ... that would work if I could remember it

OpenStudy (callisto):

This is complicated...

OpenStudy (turingtest):

I got it :) so are you with me up to the idea that tanx>x in (0,pi/2) would prove that this is decreasing?

OpenStudy (callisto):

Yes!

OpenStudy (turingtest):

okay, we can prove this using the mean value theorem...

OpenStudy (turingtest):

the mean value theorem states that f(x)=f(a)+f'(c)(x-a) for some a<c<x our f(x)=tanx, and our interval is (0,pi/2) so a=0

OpenStudy (turingtest):

getting what we need from the formula f(0)=0, f'(c)=sec^2(c) where 0<c<x<pi/2

OpenStudy (turingtest):

what do you get then by using this in the formula?

OpenStudy (callisto):

f(x)=f(a)+f'(c)(x-a) tanx = 0 + sec^2 (c) (x) tanx = [sec^2 (c)] x Doesn't look good.

OpenStudy (turingtest):

it is though, because what do we know about sec(c) for 0<c<pi/2 ???

OpenStudy (callisto):

0<c<pi/2 1<sec (c)< undefined (+infinity?!)

OpenStudy (turingtest):

yeah, but the point is that sec(c)>1 for that interval

OpenStudy (turingtest):

so tanx=sec^2(c)x 0<c<pi/2 implies that...?

OpenStudy (callisto):

1< sec^2(c) x < sec^2 (c) x x< tanx Wow!!

OpenStudy (turingtest):

tadah! :) so that completes your proof

OpenStudy (callisto):

Uh-huh! Why did you think of mean value theorem for this question????

OpenStudy (turingtest):

It's the only way I know to prove tanx>x in (0,pi/2) there may be other ways

OpenStudy (experimentx):

Given the interval is \( (0, {\pi \over 2}) \) to find the maximum or minimum, we have \[ x = \tan x\] the first part is a line y=x passing through origin while the latter one is y=tan(x)  we can see that tan(x) > x since tan(x) > slope is sec^2(x) > slope of (x) = 1 so next time tan(x) meets x in the interval pi/2, pi Hence we do not have any critical point in the interval (0, pi/2)